Shhh:

i need a in mah maths project and the topic VERIFICATION OF ALGEBRAIC IDENTITY(a+b+c)^2.

Shhh:

1 attachment
XxXNessalulbaddieXxX:

i cant even see the thing

Shhh:

i didn't did mah project so i have do thats why im noh sleeping

surjithayer:



surjithayer:

\[(a+b+c)^2=a^2+ab+ac+ab+b^2+bc+ac+bc+c^2\] \[=a^2+b^2+c^2+2ab+2bc+2ac\] mathematically \[(a+b+c)^2=a(a+b+c)+b(a+b+c)+c(a+b+c)\] \[=a^2+ab+ac+ab+b^2+bc+ac+bc+c^2\] \[=a^2+b^2+c^2+2ab+2bc+2ac\]

Shhh:

@surjithayer wrote:
\[(a+b+c)^2=a^2+ab+ac+ab+b^2+bc+ac+bc+c^2\] \[=a^2+b^2+c^2+2ab+2bc+2ac\] mathematically \[(a+b+c)^2=a(a+b+c)+b(a+b+c)+c(a+b+c)\] \[=a^2+ab+ac+ab+b^2+bc+ac+bc+c^2\] \[=a^2+b^2+c^2+2ab+2bc+2ac\]
thnx